Prove or disprove that 2 + (11)^1/2∈Q .
1
Expert's answer
2020-03-12T12:02:30-0400

"\sqrt{11}= 3.31662479036..." , which cannot be represented in the form p/q , as the fraction part is non-terminating non-recurring .

As this cannot be represented in the form of a rational number, "\sqrt{11}" is irrational.

Thus, "2+ \sqrt{11}" is also irrational.

"\implies 2+\sqrt{11} \notin Q"

Proved.